The student starts a new project making necklaces and bracelets on August 1. The student finishes a necklace every 6 days and a bracelet every 8 days.
B. What is the first date on which the student will finish both a necklace and a bracelet?
The student starts making pairs of earrings as well. Every pair of earrings takes the same number of days to complete. The student starts making pairs of earrings on a day when the student starts making both a necklace and a bracelet. The student continues to make necklaces and bracelets at the same rate as in question B. Over the next 96 days, the student makes a combined total of 36 necklaces, bracelets, and pairs of earrings.
C. How many days does it take the student to complete each pair of earrings. Show or explain your work.

Answer by greenestamps(13208)



B. August 1, plus the number of days that is the least common multiple of 6 and 8, which is 24. So 24 days after August 1.

ANSWER: August 25.

96 days means 96/6 = 16 necklaces and 96/8 = 12 bracelets; that means 36-(16+12) = 36-28 = 8 earrings.

C. 8 earrings in 96 days means 1 earring every 96/8 = 12 days.

ANSWER: 12 days